OpenStudy (bruno102):

A ladder is resting against a wall. The ladder and the ground make an angle of 35° and the ladder is 8 ft from the wall. How long is the ladder? Round your answer to the nearest hundredth.

OpenStudy (tkhunny):

I'm guessing we are to assume the wall is perpendicular to the ground. Have you considered the Tangent function?

OpenStudy (bruno102):

would it be 9.77 ft?

OpenStudy (tkhunny):

You tell me. How did you get that?

OpenStudy (tkhunny):

\(\tan(35º) = 0.7002\) \(9.77/8 = 1.22125\) Doesn't seem right.

OpenStudy (bruno102):

I don't believe I am using tangent, I believe it is cosine.

OpenStudy (tkhunny):

That would be a bad idea, since you do not know the hypotenuse length.

OpenStudy (bruno102):

But you can not use tangent in this problem because the opposite side is the side at question? Can you not do Cos(35=8/L and then do L=8/Cos(35) which is L=8/.819 and that equals L=9.768 which rounds to 9.77?

OpenStudy (tkhunny):

"The opposite side is the side at question" That is why you MUST use the TANGENT. \(\tan(35º) = \dfrac{Height}{8\;ft}\) \(Height = 8\;ft * tan(35º) = 8\;ft * 0.7002 = 5.602\;ft\) Generally, greater than 45º, the wall distance will be greater than the ground distance. Less than 45º, the wall distance will be less than the ground distance.

OpenStudy (bruno102):

Excuse me, I said that wrong, the length of the ladder is the hypotenuse.

OpenStudy (tkhunny):

It does help to write the correct question. Why do you doubt? Move on to the next one.

OpenStudy (bruno102):

The question was correct.

OpenStudy (tkhunny):

What, your ladders don't bend at right angles and crawl up next to the wall? VERY good work arguing with me and being correct.

OpenStudy (bruno102):

Thank you! I'm glad I was correct. I apologize for messing up what I was trying to say and not being clear in my reply. Sometimes it is hard to explain this kind of stuff. But anyways thanks!